40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 8 20:11:54 2012, Jenne, Update, Alignment, AS steering moved out of POY's way, 2 green beams onto PSL table AS_8Nov2012.pngREFL_8Nov2012.png
    Reply  Thu Nov 8 20:54:08 2012, Manasa, Ayaka, Update, Alignment, Reconfirming on IPPOS, IPANG and oplevs centering 
    Reply  Fri Nov 9 08:11:57 2012, Steve, Update, Alignment, nothing moved - all good opsingGOOD.pngop8h.png8amF.pngkeepfailing.png
       Reply  Fri Nov 9 11:38:38 2012, jamie, Update, General, we're closing up 
          Reply  Fri Nov 9 17:15:05 2012, Manasa, Steve, Ayaka, Update, General, We're closed! Pumping down monday morning 
Message ID: 7689     Entry time: Thu Nov 8 20:11:54 2012     Reply to this: 7690   7692
Author: Jenne 
Type: Update 
Category: Alignment 
Subject: AS steering moved out of POY's way, 2 green beams onto PSL table 

[Jenne, Jamie, Manasa, Ayaka]

 

Flipped mount of OM2, moved OM2 behind POY pickoff so we're out of the way of POY.  Adjusted and recovered rest of AS path.

We found that IPANG was not on its photodiode, but determined that it was centered on all of the in-vac mirrors, and that it was just a little bit of steering on the ETMY end out-of-vac table that needed to be done.

Got green flashes in Yarm, moved down periscope to the north by ~1 inch in order to get y green out to PSL table.  This also involved moving the steering mirror on the IOO table immediately after the down periscope to match.  We measured the MC spot positions before and after touching the periscope, and there was no significant change. 

Aligned X green to X arm (centered on ITMX, ETMX, although no flashes since we didn't move ETMX's biases around), then made sure it was centered on all of its steering mirrors, and came out of the vacuum.

Manasa took photos of all test mass chambers and the BS chamber, so we can keep up-to-date CAD drawings. 

Oplevs and IPPOS/IPANG are being centered as I type.  Manasa and Ayaka are moving the lens in front of IPANG such that we have a slightly larger beam on the QPD.

 

In the morning, Jamie is going to put apertures back on 2 of the suspended mirrors for one last check that moving things on the IOO table didn't do anything bad, but since the AS and REFL beams on those cameras didn't move significantly, we think things are fine. 

Heavy doors go on in the morning, and access connector at ~1pm, if not before lunch.  Then Steve will start pumping early Monday morning!  Hooray!

 

PS, for reference,

AS: AS_8Nov2012.png

REFL: REFL_8Nov2012.png

ELOG V3.1.3-